### v2

# Exploring New Physics with PandaX-4T Low Energy Electronic Recoil Data Xinning Zeng

Exploring New Physics with PandaX-4T Low Energy Electronic Recoil Data
Xinning Zeng
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Zihao Bo
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Wei Chen
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xun Chen
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yunhua Chen
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Zhaokan Cheng
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Xiangyi Cui
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Yingjie Fan
Department of Physics, Yantai University, Yantai 264005, China
Deqing Fang
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Zhixing Gao
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Lisheng Geng
School of Physics, Beihang University, Beijing 102206, China
Peng Huanwu Collaborative Center for Research and Education, Beihang University, Beijing 100191, China
International Research Center for Nuclei and Particles in the Cosmos & Beijing Key Laboratory of Advanced Nuclear Materials and Physics, Beihang University, Beijing 100191, China
Southern Center for Nuclear-Science Theory (SCNT), Institute of Modern Physics, Chinese Academy of Sciences, Huizhou 516000, China
Karl Giboni
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Xunan Guo
School of Physics, Beihang University, Beijing 102206, China
Xuyuan Guo
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Zichao Guo
School of Physics, Beihang University, Beijing 102206, China
Chencheng Han
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Ke Han
ke.han@sjtu.edu.cnSchool of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Changda He
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinrong He
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Di Huang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Houqi Huang
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai, 200240, China
Junting Huang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Ruquan Hou
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yu Hou
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Xiangdong Ji
Department of Physics, University of Maryland, College Park, Maryland 20742, USA
Xiangpan Ji
School of Physics, Nankai University, Tianjin 300071, China
Yonglin Ju
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Chenxiang Li
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jiafu Li
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Mingchuan Li
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Shuaijie Li
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Tao Li
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Zhiyuan Li
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Qing Lin
State Key Laboratory of Particle Detection and Electronics, University of Science and Technology of China, Hefei 230026, China
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
Jianglai Liu
jianglai.liu@sjtu.edu.cnNew Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Congcong Lu
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Xiaoying Lu
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Lingyin Luo
School of Physics, Peking University, Beijing 100871, China
Yunyang Luo
Department of Modern Physics, University of Science and Technology of China, Hefei 230026, China
Wenbo Ma
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Yugang Ma
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Yajun Mao
School of Physics, Peking University, Beijing 100871, China
Yue Meng
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Xuyang Ning
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Binyu Pang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Ningchun Qi
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Zhicheng Qian
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xiangxiang Ren
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Dong Shan
School of Physics, Nankai University, Tianjin 300071, China
Xiaofeng Shang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xiyuan Shao
School of Physics, Nankai University, Tianjin 300071, China
Guofang Shen
School of Physics, Beihang University, Beijing 102206, China
Manbin Shen
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Wenliang Sun
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yi Tao
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Anqing Wang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Guanbo Wang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Hao Wang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jiamin Wang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Lei Wang
College of Nuclear Technology and Automation Engineering, Chengdu University of Technology, Chengdu 610059, China
Meng Wang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Qiuhong Wang
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Shaobo Wang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai, 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Siguang Wang
School of Physics, Peking University, Beijing 100871, China
Wei Wang
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Xiuli Wang
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Xu Wang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Zhou Wang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yuehuan Wei
Sino-French Institute of Nuclear Engineering and Technology, Sun Yat-Sen University, Zhuhai, 519082, China
Weihao Wu
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yuan Wu
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Mengjiao Xiao
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Xiang Xiao
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Kaizhi Xiong
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Yifan Xu
School of Mechanical Engineering, Shanghai Jiao Tong University, Shanghai 200240, China
Shunyu Yao
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai, 200240, China
Binbin Yan
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Xiyu Yan
School of Physics and Astronomy, Sun Yat-Sen University, Zhuhai 519082, China
Yong Yang
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Peihua Ye
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Chunxu Yu
School of Physics, Nankai University, Tianjin 300071, China
Ying Yuan
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Zhe Yuan
Key Laboratory of Nuclear Physics and Ion-beam Application (MOE), Institute of Modern Physics, Fudan University, Shanghai 200433, China
Youhui Yun
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Minzhen Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Peng Zhang
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Shibo Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Shu Zhang
School of Physics, Sun Yat-Sen University, Guangzhou 510275, China
Tao Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Wei Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Yang Zhang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Yingxin Zhang
Research Center for Particle Science and Technology, Institute of Frontier and Interdisciplinary Science, Shandong University, Qingdao 266237, Shandong, China
Key Laboratory of Particle Physics and Particle Irradiation of Ministry of Education, Shandong University, Qingdao 266237, Shandong, China
Yuanyuan Zhang
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Li Zhao
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Jifang Zhou
Yalong River Hydropower Development Company, Ltd., 288 Shuanglin Road, Chengdu 610051, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Jiaxu Zhou
SJTU Paris Elite Institute of Technology, Shanghai Jiao Tong University, Shanghai, 200240, China
Jiayi Zhou
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
Ning Zhou
New Cornerstone Science Laboratory, Tsung-Dao Lee Institute, Shanghai Jiao Tong University, Shanghai 201210, China
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Shanghai Jiao Tong University Sichuan Research Institute, Chengdu 610213, China
Jinping Deep Underground Frontier Science and Dark Matter Key Laboratory of Sichuan Province
Xiaopeng Zhou
zhou_xp@buaa.edu.cnSchool of Physics, Beihang University, Beijing 102206, China
Yubo Zhou
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
Zhizhen Zhou
School of Physics and Astronomy, Shanghai Jiao Tong University, Key Laboratory for Particle Astrophysics and Cosmology (MoE), Shanghai Key Laboratory for Particle Physics and Cosmology, Shanghai 200240, China
(February 7, 2025)
Abstract
New particles beyond the standard model of particle physics, such as axions, can be effectively searched through their interactions with electrons.
We use the large liquid xenon detector PandaX-4T to search for novel electronic recoil signals induced by solar axions, neutrinos with anomalous magnetic moment, axionlike particles, dark photons, and light fermionic dark matter.
A detailed background model is established using the latest datasets with 1.54 exposure.
No significant excess above the background has been observed, and we have obtained competitive constraints for axion couplings, neutrino magnetic moment, and fermionic dark matter interactions.
Large liquid xenon (LXe) detectors located deep underground have unique advantages in searching for rare signals from dark matter (DM) and neutrino-related physics [1, 2, 3, 4, 5].
In addition to the traditional DM-induced nuclear recoil (NR) events, particles from many new physics models can produce novel electronic recoil signals (NERS), which have attracted increasing attention recently [6, 7, 8, 9].
Axions [10, 11, 12] can be produced in the Sun via the ABC effect (atomic recombination and deexcitation, bremsstrahlung, and Compton scattering processes) [13, 14], the Primakoff conversion [15], and the nuclear magnetic transitions [16], related to the axion-electron coupling , axion-photon coupling , and effective axion-nucleon coupling , respectively.
The corresponding effective Lagrangian is given by [17]:
where is the axion field, the electron, the photon, and the nucleon isospin doublet .
and are model-dependent isoscalar and isovector axion-nucleon couplings and is the Pauli matrix.
The Sun also produces a vast number of neutrinos, which may carry anomalous magnetic moment due to the Majorana nature of neutrino or exotic new physics beyond the standard model (SM) [18].
With their kinetic energy governed by the Sun’s temperature, these particles can produce electronic recoil (ER) events with typical energy from keV to tens of keV.
Within this energy range, alternative light dark matter models such as axionlike particles (ALPs), dark photons [19], or light fermionic DM converting into active neutrinos (DM-) [20, 21] can also manifest themselves with characteristic ER spectra.
ALPs interact with SM particles similarly to axions, while dark photons interact through kinetic mixing with SM photons.
The vector and axial-vector interactions among the dimension-six effective operators of DM- conversion can be probed experimentally. Considering the atomic effect, the differential cross section with respect to the ionized electron energy is
where and are the principal quantum number and angular momentum.
— is the energy of outgoing neutrino and is the energy of the accompanying x-ray, which has the same energy as the binding energy of state .
is the particle scattering amplitude and
is the atomic factor [22].
The masses of electron and DM particle are denoted as and , respectively.
This Letter reports an updated NERS search using the low-energy ER data below 30 keV from PandaX-4T, an LXe detector operating in the China Jinping Underground Laboratory (CJPL), with a total exposure of 1.54 tonyr.
In different NERS analyses, the same method of searching for spectral features on top of the expected background is used.
We leave detailed production and detection cross sections and expected NERS spectra in the Supplemental Material [23] for brevity.
To set the stage, the signals of ABC solar axions and the DM- conversion of keV/ are shown in Fig. 1, with spectra at the low and high end of the region of interest (ROI).
The keys to this analysis lie in the understanding of the detector responses, as well as robust prior knowledge of the background from data outside the ROI.
Figure 1: Theoretical (shaded) and expected energy spectra (solid for Run0, dashed for Run1) for ABC axion with a coupling =3.1 (red shaded for theoretical, yellow unshaded for expected) and DM- conversion with = 170 keV and cm2 (blue shaded for theoretical, purple unshaded for expected).
The solid and dashed green lines depict the signal detection efficiencies for Run0 and Run1, including reconstruction, quality cuts, and ROI, with their respective pink and green bands indicating the uncertainties.
Table 1: Operating conditions of Run0 and Run1. Duration represents the calendar days, and live time refers to the data-taking time after removing low-quality data periods. is the average electron lifetime of each dataset. and are voltages of the cathode and gate. and values are determined by likelihood fits to low energy calibration data [41].
Run
Run0
Run1
Set
1
2
3
4
5
1
2
3
4
5
6
Duration (days)
1.96
13.54
5.53
37.22
36.61
9.41
10.50
6.18
14.00
38.44
85.08
Live Time (days)
1.82
12.30
2.14
33.01
34.29
8.07
8.53
5.56
12.19
34.45
77.86
807.2
958.1
812.5
1110.4
1275.4
631.2
642.3
975.7
831.5
1171.7
1719.6
(-kV)
20
18.6
18
16
16
16
(-kV)
4.9
4.9
5
5
5
6
(PE/)
The PandaX-4T utilizes a time projection chamber (TPC), with 3.7 ton of LXe in the sensitive volume, to measure the precise position, timing, and energy of each event [42].
Energy release in xenon produces prompt scintillation photon signals () and ionized electrons.
Electrons drift upward under the influence of a downward electric field in the sensitive volume, and the delayed electroluminescence signals () are generated in the gaseous xenon region.
Photons from both signals are collected by arrays of Hamamatsu R11410-23 three-inch photomultiplier tubes (PMTs) at the top and bottom of the TPC.
Combining the photon pattern on the PMT arrays and the time delay between and , the three-dimensional vertex of a given event can be deduced.
The electron-equivalent recoil energy is reconstructed as
in which () is the corrected charge of ( collected by the bottom PMTs), and eV is the mean energy to produce a single photon or ionized electron in liquid xenon.
Using avoids the influence of potential saturation and dead channel effects from the top PMTs.
is the photon detection efficiency for .
is the product of electron extraction efficiency and the single-electron gain of .
Both parameters are determined via calibrations [41, 1].
For this analysis, we utilize the data from Run0 (Nov. 28, 2020 - Apr. 16, 2021) and Run1 (Nov. 16, 2021 — May 15, 2022), which are divided into subsets based on different electron lifetimes, liquid levels, cathode voltages, and gate voltages, as summarized in Table 1.
The fiducial masses optimized to reject external backgrounds from detector materials are and ton for Run0 and Run1, respectively.
The signal detection efficiency for Run0 (Run1), including reconstruction efficiency, quality cuts efficiency, and ROI efficiency, is estimated by both the data-driven method and waveform-simulation method [41, 43].
The low threshold of the ROI is 2 photoelectrons (PE) for , and the upper limit is 30 keV in .
A 99.5% ER acceptance cut in the space of log vs is applied, which can reject more than 66% (99%) of the NR events from neutrons (the coherent elastic scattering of solar 8B neutrino on xenon nucleus) and 95% of NR-like background originating from radon plate-out from the polytetrafluoroethylene reflector surface in the TPC.
The overall efficiency is estimated to be () for Run0 (Run1), and the energy dependence is shown in Fig. 1.
Above 7 keV, the efficiency plateaus and the average value is () for Run0 (Run1).
Major background sources of the PandaX-4T experiment can be categorized into three types.
Environmental backgrounds are from cosmic muons, solar neutrinos, and radioisotopes from surrounding rocks in the lab.
External backgrounds originate from detector components such as PMTs.
Internal backgrounds refer to radioisotopes of xenon and other radio impurities dissolved in the target.
Table 2: Summary of background expectations and best fit of the background-only hypothesis of each component for Run0 and Run1. Systematic uncertainties from the signal model and detection efficiencies are included in the uncertainties presented.
Run
Run0
Run1
Expected
Fitted
Expected
Fitted
Tritium
578.6 32.5
214Pb
327.2 18.8
328.0 17.1
212Pb
57.8 14.7
85Kr
94.2 47.3
Material
49.4 3.3
136Xe
36.9 2.5
127Xe
6.1 0.3
124Xe
2.3 0.6
Solar
43.0 4.6
Accidental
7.6 2.4
Total
1196.632.6
Observed
1197
1431
With 2.4 km rock overburden of the Jinping Mountain, the cosmic muon flux at CJPL is measured to be [44], leading to a negligible muon-induced background in the ROI.
Solar neutrinos produced via proton-proton fusion and 7Be electron capture are calculated using standard solar model, three-flavor neutrino oscillation, standard model neutrino-electron scattering, as well as updated treatment of the xenon atomic effects [45, 46, 47].
A 10% uncertainty is used based on Borexino measurements [48].
The concentration of K, Th, and U of the CJPL lab environment are measured at Bq/kg level [49].
With at least 4 m of water shielding in each direction [42], the contribution from surrounding rocks is suppressed to a negligible level.
The concentrations of 232Th and 238U in ultra-pure water shielding itself are measured to be ppt (part per trillion) and are expected to contribute events/ton/yr/keV [50] in the ROI.
The background from rays of detector materials is calculated based on the ER spectrum in the MeV energy range, where characteristic full-absorption peaks can help constrain the radioactivities of different components in situ [51].
The external backgrounds contribute a flat shape in the ROI according to a custom geant4-based simulation package BambooMC [52, 53], with the dominant contributions from the stainless steel vessels and PMTs.
The expected background counts in the ROI are and events for Run0 and Run1, extrapolated from the results of Ref. [51] with the help of BambooMC.
Internal backgrounds in the ROI are mainly from decays of the progenies of radon, especially 222Rn dissolved in LXe.
The decay chain,
practically stops at the long-lived 210Pb.
The concentration of 222Rn is () Bq/kg for Run0 (Run1), determined from the number of decay events.
The decay rate of 214Pb cannot be accurately determined from 222Rn.
A depletion effect, i.e., decreased concentration of longer-lived daughter nuclide compared to the parent in LXe, has been reported [54].
The concentration of 214Pb for Run0 is Bq/kg, derived from the spectral fit in 200-1000 keV [55].
For Run1, the concentration is estimated to be Bq/kg, based on the ratio of 222Rn concentrations relative to Run0.
For 220Rn decay chain, the dominating background comes from 212Pb, which is estimated to be Bq/kg with the depletion effect taken into account, using the level of its daughter 212Po decays as the proxy.
Another emission source is the 85Kr.
The krypton concentration in Run0 (Run1) is 0.52 0.27 (0.94 0.28) ppt, estimated with the coincidence of 85Kr decay to metastable state 85mRb and the subsequent deexcitation to ground state 85Rb.
The predominant contribution to the uncertainty is the limited statistics of - coincidence, which are 4 events in Run0 and 12 events in Run1.
Assuming an isotopic abundance of , 94.2 (308.1) 85Kr events are in our ROI for Run0 (Run1).
In the Run0 data, tritium events are observed at a concentration of mol/mol [56].
The contamination is likely introduced to xenon during the PandaX-II calibration runs [6].
After a dedicated tritium removal campaign after Run0, the tritium contribution in Run1 drops to mol/mol.
We do not impose any constraint of tritium in our spectral fit.
Background contributions from xenon isotopes are considered, notably the 136Xe and 124Xe, along with cosmogenic 127Xe.
The isotopic abundance of 136Xe and 124Xe are measured to be 8.7% and 0.1%.
The half-life of 136Xe’s double beta decay is measured in Ref. [51], leading to an estimated background contribution of events for Run0 and events for Run1.
Double electron capture events of LL-shell of 124Xe with an expected energy of approximately 10 keV are estimated based on the theoretical branching ratio and the measured 124Xe half-life from XENON1T [57].
The expected number of events for Run0 and Run1 is and , respectively.
The decay from -shell electron capture (5.2 keV) of 127Xe also contributes to our ROI, which can be determined by the 408 keV peak (-shell capture mixed with the 375 keV -ray from 127I).
In addition to the physical background, accidental coincidence (AC) events resulting from the random pairing of isolated and signals also contribute to the background.
These spurious events are estimated to be and in Run0 and Run1, respectively [58].
The background levels do not scale with the exposure due to different cut efficiencies in the two runs.
Figure 2: Joint fit of the Run0 (top) and Run1 (bottom) ER data with background-only hypothesis.
Backgrounds are stacked in the sequence presented by the legend.
The 1 and 2 residual bands are depicted with dark cyan and yellow bands, respectively.
Details of the fit results are listed in Table 2.
Our final analysis immediately followed the unblinding of Run0 and Run1 data in Ref. [1], and no post-unblinding adjustments were made on the event selection cuts and background models.
In total, 1197 (1431) ER events are selected in Run0 (Run1), different from those in Ref. [1] only due to a broader range of ROI, and the removal of NR-like data beyond the 99.5% ER acceptance.
A binned one-dimensional likelihood fit in energy based on HistFitter [59, 60] is performed to test the background-only hypotheses.
Background estimates in Table 2 are taken with prior uncertainties as constraints.
Systematics, including those corresponding to the signal model in PandaX-4T [41] and detection efficiency (see Fig. 1), are also incorporated into the fit as nuisance parameters.
Among all signal model uncertainties, the electron recombination rate (affecting the energy division between scintillation and ionization) impacts the low energy spectrum, as a change in the light yield mostly influences events close to the lower threshold of PE [9].
The propagated uncertainty on the spectrum is within 3.7% below 2 keV and has no impact above.
The best fit of a background-only model is shown in Fig. 2, with the number of fitted background events summarized in Table 2.
The fitted background contributions agree with the expected values, with all fitted nuisance parameters within of the input values.
The ROI’s average fitted ER background rate (including efficiency) is and events/ton/yr/keV in Run0 and Run1, respectively.
The data are consistent with the background-only model with a value of 0.16.
Figure 3: The 90% CL upper limits of couplings of different NERS. (a) for ABC axion; (b) for 57Fe axion; (c) for Primakoff axion; (d) for solar neutrino with anomalous magnetic moment ( is the Bohr magneton); (e) for ALPs; (f) mixing angle for dark photons; (g)(h) cross sections of light fermionic DM- conversion for the vector and axial-vector interactions. In (g)(h), the limits (blue solid line) and sensitivity median (blue dashed line) of Run0 published results [7], denoted as PandaX-4T (2022), are also overlaid.
In scenarios (e) to (h), we assume that each type of dark matter candidate constitutes 100% of the local dark matter.
The limits from other experiments [61, 62, 63, 64, 65, 17, 66, 67, 48, 68, 69, 70, 71, 72, 73, 8, 9] are also shown.The shaded orange region in (a)(c) represents predictions derived from two benchmark QCD axion models: KSVZ [74, 75] and DFSZ [76, 77].
NERS hypotheses are tested with the background-plus-signal fits using the Run0 and Run1 data.
The best fit of the ABC axions hypothesis yields events (ranged to positive only), while the best fits for axions originating from the Primakoff conversion and 57Fe are null, assuming an axion mass of keV/.
For the anomalous neutrino magnetic moment hypothesis, the best fit is events, also compatible with zero.
ALP signals, as well as those from dark photons and DM- conversion, exhibit a characteristic peak structure that correlates with their mass.
The mass scanning for ALPs and dark photons is conducted within the range of 1 to 25 keV/, while for DM- conversion, the range is from 10 to 180 keV/.
The data exhibit an upward fluctuation in the 3 - 5 keV range for both Run0 and Run1, which results in a local significance of 1.9 for ALPs and dark photons at a mass of 3.3 keV/, and 1.6 for light DM at a mass of 60 keV/ (peak of recoil energy at 4.1 keV) relative to the background-only hypothesis [78].
As none of the local excess is beyond 3, according to the conventions in Ref. [79], we proceed with the upper limit interpretation of our data using the profile likelihood ratio approach outlined in Ref. [78].
The upper limits with 90% confidence level (CL) of axion couplings (), , kinetic mixing of dark-to-SM photon and DM- conversion cross sections () are shown in Fig. 3.
The corresponding 1 sensitivities are shown as green bands in the figure.
In the scanning over DM mass in the last four panels in Fig. 3, local data fluctuations cause corresponding variations in the derived limits.
For conservativeness, limits due to upward fluctuation of data are reported as they are, but those due to downward fluctuation are power constrained to -1 boundary of the sensitivity band [80].
For comparison, constraints from other direct detection experiments and astrophysical observations are overlaid.
The upper limit derived from ABC axions is for keV/.
For axions produced by the 14.4 keV 57Fe transition, are constrained to less than for keV/, which is as competitive as the constraints from XENONnT [8].
The axion-photon coupling is constrained to GeV-1 for keV/, which represents one of the most competitive limits on from a terrestrial experiment, and disfavoring KSVZ [74, 75] and DFSZ [76, 77] axions above a few eV/.
The upper limit on the anomalous neutrino magnetic moment is .
Compared to our previous results [6], the constraints on and have been improved by about 2 times.
Our results for ALPs and dark photons limits are slightly weaker than those from XENONnT [8].
Our sensitivities for DM- conversion are notably improved compared to the previous best results in Ref. [7], except at around 60 keV/ and between 150 - 180 keV/ due to the upward data fluctuation at around 3 keV and 24 - 25 keV.
In summary, we have searched for NERS, including solar axions, neutrinos with anomalous magnetic moment, ALPs, dark photons, and light fermionic DM- conversion, with the 1.54 of PandaX-4T low-energy electronic recoil data.
No statistically significant NERS have been observed.
Consequently, stringent upper limits are established on NERS.
In particular, we report leading limits on solar 57Fe axions and light fermionic DM- conversion, the latter of which is the most stringent result for most of the mass range of 10 - 180 keV/.
After a long shutdown due to the construction work in CJPL, PandaX-4T has restarted detector commissioning.
Improved search sensitivities for NERS are expected with new data.
This project is supported in part by grants from National Key R&D Program of China (No. 2023YFA1606200, No. 2023YFA1606201, National Science Foundation of China (No. 12090060, No. 12090061, No. 12105008, No. U23B2070), and by Office of Science and Technology, Shanghai Municipal Government (Grant No. 21TQ1400218, No. 22JC1410100, No. 23JC1410200, No. ZJ2023-ZD-003). We thank the support by the Fundamental Research Funds for the Central Universities, Hongwen Foundation in Hong Kong, New Cornerstone Science Foundation, Tencent Foundation in China, and Yangyang Development Fund. Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd. for indispensable logistical support and other help.
Viaux et al. [2013]N. Viaux, M. Catelan, P. B. Stetson, G. G. Raffelt, J. Redondo, A. A. R. Valcarce, and A. Weiss, Phys. Rev. Lett. 111, 231301 (2013).